# Graph Attention Network-based Block Propagation with Optimal AoI and Reputation in Web 3.0

### v1

@@ IV-B3 Final reputation for subjective logic
For the calculation of final miner reputations, it is important to reasonably consider the weights of local reputation opinions and recommended reputation opinions, thereby ensuring the final reputation of miners is more accurate and trustworthy. Based on the above analysis, the final reputation opinion of miner $i$ to miner $j$ is composed of a tuple vector $\Phi_{i\to j}^{final}:=\{T_{i\to j}^{final},F_{i\to j}^{final},U_{i\to j}^{final}\}$, which is given by
@@
IV-C Problem Formulation
@@
$\begin{split}&\min_{\bm{\pi}}\>\sum_{i=2}^{m}\Delta_{i}=\Gamma_{i}+\frac{1}{\mu}+\frac{\mu\Gamma_{i}^{3}}{1-\mu\Gamma_{i}}\\ &\>\>\mathrm{s.t.}\>R_{\pi_{l}\rightarrow\pi_{l+1}}^{final}>\sigma,l\in\{1,\ldots,m-1\},\end{split}$ (23)
where $\sigma\in(0,1)$ is a miner reputation threshold for ensuring block propagation. Considering $\Gamma_{i}<(1/\mu)$[24], it is easy to prove that the AoI $\Delta_{i}$ will increase as block propagation time $\Gamma_{i}$ increases. Moreover, it is obvious that $\Gamma_{i}$ is a monotonic function concerning the distance $d_{i}$ between miner $(i-1)$ and miner $i$ according to (11). Motivated by the above analysis, we propose a GAT-based block propagation optimization model to solve the optimization problem, described in Section V.
V Graph Attention Network-based Block Propagation Model
In this section, we propose a GAT-based block propagation optimization model for blockchain-enabled Web 3.0, which minimizes the average AoI of block propagation between adjacent miners with the specific reputation constraint, thereby obtaining the optimal block propagation trajectory.
@@ V-A Encoder
where $\bm{W}^{h}\in\mathbb{R}^{d_{h}\times d_{x}}$ and $\bm{b}^{h}\in\mathbb{R}^{d_{h}}$. After obtaining the embedding $\bm{h}_{i}^{(0)}$, it is sent to the graph attention layer and updated with $N$ GAT layers. We denote $\bm{h}_{i}^{(r)}$ as the miner embeddings calculated by GAT layer $r\in\{0,1,\ldots,N\}$. Specifically, $\bar{\bm{{h}}}^{(graph)}$ is denoted as the graph embedding, which is the aggregated embedding of the input graph as the average of final miner embeddings, given by[23]
@@
V-A1 Multi-Head Attention (MHA) layer
The MHA layer is beneficial to learning information from different aspects than only a single-head attention mechanism[32]. Inspired by [31, 33, 34], we utilize the MHA layer to model the relevance between miners on the graph. In the MHA layer, the value111 In attention mechanisms, a query represents the focused content, a key serves as a reference, and a value corresponds to the actual information associated with the key[31]. of a miner is the compatibility of the query and the key from its neighbor[31]. We denote the number of attention heads as $y\in\{1,\ldots,Y\}$ and consider a sequence of query $\bm{Q}=\{\bm{q}_{1y}^{(r)},\ldots,\bm{q}_{iy}^{(r)},\ldots,\bm{q}_{My}^{(r)}\}$, key $\bm{K}=\{\bm{k}_{1y}^{(r)},\ldots,\bm{k}_{iy}^{(r)},\ldots,\bm{k}_{My}^{(r)}\}$, and value $\bm{V}=\{\bm{v}_{1y}^{(r)},\ldots,\bm{v}_{iy}^{(r)},\ldots,\bm{v}_{My}^{(r)}\}$[35]. For miner $i$, $\bm{q}_{iy}^{(r)}$, $\bm{k}_{iy}^{(r)}$, and $\bm{v}_{iy}^{(r)}$ are calculated by projecting the embedding $\bm{h}_{iy}^{(r-1)}$, which are given by
$\bm{q}_{iy}^{(r)}=\bm{W}^{Q}\bm{h}_{iy}^{(r-1)},$ (29)
$\bm{k}_{iy}^{(r)}=\bm{W}^{K}\bm{h}_{iy}^{(r-1)},$ (30)
$\bm{v}_{iy}^{(r)}=\bm{W}^{V}\bm{h}_{iy}^{(r-1)},$ (31)
@@ V-A1 Multi-Head Attention (MHA) layer
Based on (29), (30), and (31), we compute the compatibility $\bm{u}_{ijy}^{(r)}\in\mathbb{R}$ by combining the query $\bm{q}_{iy}^{(r)}$ with the key $\bm{k}_{jy}^{(r)}$ of miner $j$ as the dot-product function[31]
$\bm{u}_{ijy}^{(r)}=\begin{cases}\frac{\left(\bm{q}_{iy}^{(r)}\right)^{T}\bm{k}_{jy}^{(r)}}{\sqrt{d_{k}}}&\text{if miner $i$ is adjacent to miner $j$,}\\ -\infty&\text{otherwise,}\end{cases}$ (32)
where the compatibility of non-adjacent miners is considered as $\infty$- ∞, which is effective in preventing block propagation between non-adjacent miners. Then, based on the compatibilities $\bm{u}_{ijy}^{(r)}$, the attention weight $\bm{a}_{ijy}^{(r)}\in[0,1]$ is given by[23]
$\bm{a}_{ijy}^{(r)}=softmax\left(\bm{u}_{ijy}^{(r)}\right)=\frac{e^{\bm{u}_{ijy}^{(r)}}}{\sum_{j=1}^{J}e^{\bm{u}_{ijy}^{(r)}}},$ (33)
where $\{1,\ldots,j,\ldots,J\}\subset\mathcal{M}$ is the adjacent miner set of miner $i$. Then, based on (31) and (33), the result vector $\bm{h}_{iy}^{\prime}(r)$, combining $\bm{a}_{ijy}^{(r)}$ with $\bm{v}_{jy}^{(r)}$, is expressed as[23]
$\bm{h}_{iy}^{\prime}{}^{(r)}=\sum_{j=1}^{J}\bm{a}_{ijy}^{(r)}\bm{v}_{jy}^{(r)},$ (34)
@@
V-A2 Batch Normalization (BN) layer
@@
V-A3 Feed-Forward (FF) layer
@@
V-A4 Graph attention network layer
@@ V-B Decoder
Here, [,,,][$\cdot$,$\cdot$,$\cdot$,][ ⋅ , ⋅ , ⋅ , ] is the horizontal concatenation operator. For $t=1$, we introduce learnable $d_{h}$-dimensional parameters $\bm{v}^{1}$ and $\bm{v}^{2}$ as input placeholders. For $t=2$, the placeholder $\bm{v}_{1}$ would be replaced by the embedding $\bm{h}_{\pi_{1}}^{(N)}$. For $t>2$, the placeholder $\bm{v}_{2}$ and the embedding $\bm{h}_{\pi_{1}}^{(N)}$ would be replaced by the embeddings $\bm{h}_{\pi_{t-1}}^{(N)}$ and $\bm{h}_{\pi_{t-2}}^{(N)}$ of miners $\pi_{t-1}$ and $\pi_{t-2}$. Note that the result vector $\bm{h}^{(d)}$ is $(3\cdot d_{h})$-dimension to align with the miner embeddings $\bm{h}_{\pi_{t}}^{(N)}$.
@@ VI Model Training
1 for ${i}=1,\ldots,M$ do
2       Record the miner interactions $\alpha^{t_{k}}$ and $\beta^{t_{k}}$. Calculate the miner reputation $R_{i\to j}^{final}$.
3Initialize $\bm{\theta}$, $\bm{\theta}^{BL}$leftarrow $\bm{\theta}$. for $epoch=1,\ldots,E$ do
4       for $step=1,\ldots,T$ do
5             $x_{i}\leftarrow$ RandomInstance( )[23], $\forall{i}\in\{1,\ldots,B\}$. for ${i}=1,\ldots,M$ do
6                   if $R_{i\to j}^{final}$ $\sigma$ then
7                         Mask miner $i$ to avoid the low-reputation miners.
8
9            $\bm{\pi}_{i}\leftarrow$ SampleRollout($G_{i},\bm{p}_{\bm{\theta}}$)[23], $\forall{i}\in\{1,\ldots,B\}$. $\bm{\pi}_{i}^{BL}\leftarrow$ GreedyRollout($G_{i},p_{\bm{\theta}}^{BL}$)[23], $\forall{i}\in\{1,\ldots,B\}$. $\nabla\mathcal{L}\leftarrow\sum_{i=1}^{B}\left(L(\bm{\pi_{i}|G)})-L(\bm{\pi_{i}^{\mathrm{BL}}|G)})\right)\nabla_{\bm{\theta}}\log p_{\bm{\theta}}(\bm{\pi_{i}})$. $\bm{\theta}\leftarrow$ Adam($\bm{\theta},\nabla\mathcal{L}$).
10      if OneSidedPairedTTest($p_{\bm{\theta}},p_{\bm{\theta}}^{BL}$) ¡ $\alpha^{gat}$[23] then
11             $\bm{\theta}^{BL}\leftarrow\bm{\theta}$.
12
Feed the miners with their reputations into the trained network. Calculate the minimized overall AoI and the total value of miner reputations. Output: The optimal block propagation trajectory, the minimized overall AoI, and the total value of miner reputations.
Figure 3: The overall AoI of block propagation corresponding to different numbers of miners.
Figure 3: The overall AoI of block propagation corresponding to different numbers of miners.
@@
Furthermore, concerning the algorithm itself, each rollout incurs an additional forward pass, resulting in a $50\%$ increase in computational requirements[23]. However, given that the baseline policy remains constant throughout an entire epoch, we can streamline the process of sampling data and computing baselines by employing larger batch sizes. This efficiency is possible due to reduced memory requirements, allowing computations to operate in pure inference mode [23]. Note that the computational complexity of Algorithm 1 is $O(M)$[37], indicating that Algorithm 1 is efficient. Finally, after training the network, the miners, along with their reputations, are input into the trained network. The network minimizes the overall AoI of block propagation and calculates the total reputation of miners, generating the optimal block propagation trajectory for a specific number of miners as the final output.
@@ VII Numerical Results
In this section, we first compare the proposed GAT model with other conventional routing algorithms: i) Greedy algorithm, where miners consistently select the nearest adjacent miner to propagate the block; ii) Gossip algorithm, where miners randomly propagate the block to their adjacent miners. At the same time, we separately add the reputation constraints to the three proposed algorithms. Then, we present the optimal block propagation trajectories under different miner numbers in the proposed GAT model. Finally, we evaluate the impact of baselines and learning rate settings on the performance of the proposed GAT model.
@@
For the parameter settings of the experiment, the parameters $\bm{\theta}$ of the GAT network are initialized uniformly within the range $\big{(}-1/\sqrt{d_{f}},1/\sqrt{d_{f}}\big{)}$, where $d_{f}=2$ represents the input dimension of miners. Moreover, we use $N=3$ layers in the encoder, which is a thoughtful balance between the quality of results and computational complexity. We set $r^{ratio}=3/4$ to ensure that the block propagates to miners over $50\%$ for validation. Additionally, the training spans $100$ epochs using dynamically generated training data, processing $2500$ batches of $512$ instances per epoch. Note that we run the experiments on NVIDIA GeForce RTX 3080Ti, and the main parameters of this paper are listed in Table II[38, 39, 40, 25].
Figure 4 shows the overall AoI of block propagation corresponding to different numbers of miners utilizing different routing algorithms. For convenience, we denote ”GAT with Reputation” to represent the GAT model with the reputation mechanism, which is the same as the ”Greedy with Reputation” and ”Gossip with Reputation”. From Fig. 4, we can see that both the proposed GAT model and the GAT model with the reputation mechanism consistently outperform the other four algorithms. When $M=49$, the advantage is particularly pronounced, where the AoI of the GAT model is $2.5\%$ lower than that of the Greedy algorithm and $24.4\%$ lower than that of the Gossip algorithm, while the AoI of the GAT model with the reputation mechanism is $2.4\%$ lower than the Greedy algorithm with the reputation mechanism and $22.6\%$ lower than that of the Gossip algorithm with the reputation mechanism. The reason is that the GAT model focuses on minimizing the overall AoI of block propagation by considering the global structure of the miner network, which can obtain the optimal block propagation trajectory, thereby achieving block propagation optimization. In contrast, the Greedy algorithm only focuses on the selection of the current step and ignores the global structure, while the Gossip algorithm randomly selects miners without optimizing AoI, resulting in a much lower effect than other algorithms.
Figure 4 depicts the total value of miner reputations corresponding to different numbers of miners under different routing algorithms, displaying a monotonic rise with an increase in the number of miners. As we can see, the algorithms with the miner reputation mechanism can obtain the propagation trajectory with a high total of value miner reputations, thus effectively ensuring reliability in block propagation. Moreover, the GAT model with the reputation mechanism performs best among other algorithms and the reputation gap becomes more obvious with a larger number of miners. For example, comparing the Greedy algorithm with the GAT model with the reputation mechanism, the reputation of the GAT model with the reputation mechanism is $7.4\%$ higher than that of the Greedy algorithm at $M=9$, which outstandingly increases to $34.77\%$ at $M=49$. Combining Fig. 4 and Fig. 4, we can conclude that algorithms with the reputation mechanism inevitably improve the reliability of block propagation at the cost of increasing AoI [10]. However, the improvement in the reliability of block propagation with a relatively small increase in AoI is reasonable. Overall, the GAT model with the reputation mechanism stands out as the most exceptional performer among other algorithms.
@@
In Fig. 5, we present the optimal block propagation trajectory corresponding to different miner numbers $M=\{9,19,29,39,49\}$. For better distinction, we draw yellow points to denote the low-reputation miners and green points to denote the high-reputation miners, where the reputation values are calculated by the proposed reputation mechanism[30]. Moreover, the blue arrows point to the optimal block propagation trajectories generated by the proposed GAT model. We can observe that the proposed GAT model can obtain optimal block propagation trajectories based on different numbers of miners, which are clearly organized and have no unreasonable costs in AoI for efficiency. At the same time, the block propagation trajectories can perfectly prevent the blocks from propagating to the low-reputation miners, which ensures the reliability of block propagation.
@@
Figure 7 depicts the network loss corresponding to training epochs with four learning rate schemes for miner numbers $M=19$ and $M=49$. As shown in Fig. 7(a), the GAT model with the learning rate $lr=10^{-3}$ outperforms the GAT model with the learning rate $lr=10^{-4}$, with marginal improvement from learning rate decay[23]. However, in Fig. 7(b), the loss of the GAT model with the learning rate $lr=10^{-3}$ converges slowly and fails to fully converge after $100$ epochs with a larger loss. Conversely, employing a learning rate of $lr=10^{-3}$ with a decay of $0.96^{\text{epoch}}$ yields better model performance, and further improvement may be achievable with prolonged training. Furthermore, motivated by the above analysis, we conclude that employing a higher learning rate of $lr=10^{-3}$ accelerates the initial learning but necessitates decay for convergence[23], especially in the case of a larger number of miners. Moreover, a smaller learning rate of $lr=10^{-4}$ shows increased stability without the need for decay, albeit with a limited impact on loss reduction.

### v2

@@ IV-B3 Final reputation for subjective logic
For calculating final miner reputations, it is important to consider the weights of local reputation opinions reasonably and recommended reputation opinions, thereby ensuring the final reputation of miners is more accurate and trustworthy. Based on the above analysis, the final reputation opinion of miner $i$ to miner $j$ is composed of a tuple vector $\Phi_{i\to j}^{final}:=\{T_{i\to j}^{final},F_{i\to j}^{final},U_{i\to j}^{final}\}$, which is given by
@@
V Graph Attention Network-based Block Propagation Model
V-A Problem Formulation
@@
$\begin{split}&\min_{\bm{\pi}}\>\sum_{i=2}^{m}\Big{(}\Delta_{i}=\Gamma_{i}+\frac{1}{\mu}+\frac{\mu\Gamma_{i}^{3}}{1-\mu\Gamma_{i}}\Big{)}\\ &\>\>\mathrm{s.t.}\>R_{\pi_{l}\rightarrow\pi_{l+1}}^{final}>\sigma,l\in\{1,\ldots,m-1\},\end{split}$ (26)
where $\sigma\in(0,1)$ is a miner reputation threshold for ensuring block propagation. Considering $\Gamma_{i}<(1/\mu)$[26], it is easy to prove that the AoB $\Delta_{i}$ will increase as block propagation time $\Gamma_{i}$ increases. Moreover, it is obvious that $\Gamma_{i}$ is a monotonic function concerning the distance $d_{i}$ between miner $(i-1)$ and miner $i$ according to (12). Motivated by the above analysis, we propose a GAT-based block propagation optimization model to solve the optimization problem, described in Section V.
In this section, we propose a GAT-based block propagation optimization model for blockchain-enabled Web 3.0, which minimizes the average AoB of block propagation between adjacent miners with the specific reputation constraint, thereby obtaining the optimal block propagation trajectory.
@@ V-B Encoder
where $\bm{W}^{h}\in\mathbb{R}^{d_{h}\times d_{x}}$ and $\bm{b}^{h}\in\mathbb{R}^{d_{h}}$. After obtaining the embedding $\bm{h}_{i}^{(0)}$, it is sent to the graph attention layer and updated with $N$ GAT layers. We denote $\bm{h}_{i}^{(r)}$ as the miner embeddings calculated by GAT layer $r\in\{1,\ldots,N\}$. Specifically, $\bar{\bm{{h}}}^{(graph)}$ is denoted as the graph embedding, which is the aggregated embedding of the input graph as the average of final miner embeddings, given by[25]
@@
V-B1 Multi-Head Attention (MHA) layer
The MHA layer is beneficial to learning information from different aspects than only a single-head attention mechanism[32]. Inspired by [31, 33, 34], we utilize the MHA layer to model the relevance between miners on the graph. In the MHA layer, the value111 In attention mechanisms, a query represents the focused content, a key serves as a reference, and a value corresponds to the actual information associated with the key[31]. of a miner is the compatibility of the query and the key from its neighbors[31]. We denote the number of attention heads as $y\in\{1,\ldots,Y\}$ and consider a sequence of query $\bm{Q}=\{\bm{q}_{1}^{(r)},\ldots,\bm{q}_{i}^{(r)},\ldots,\bm{q}_{M}^{(r)}\}$, key $\bm{K}=\{\bm{k}_{1}^{(r)},\ldots,\bm{k}_{i}^{(r)},\ldots,\bm{k}_{M}^{(r)}\}$, and value $\bm{V}=\{\bm{v}_{1}^{(r)},\ldots,\bm{v}_{i}^{(r)},\ldots,\bm{v}_{M}^{(r)}\}$[35]. For miner $i$, $\bm{q}_{i}^{(r)}$, $\bm{k}_{i}^{(r)}$, and $\bm{v}_{i}^{(r)}$ are calculated by projecting the embedding $\bm{h}_{i}^{(r-1)}$, which are given by
$\bm{q}_{i}^{(r)}=\bm{W}^{Q}\bm{h}_{i}^{(r-1)},$ (32)
$\bm{k}_{i}^{(r)}=\bm{W}^{K}\bm{h}_{i}^{(r-1)},$ (33)
$\bm{v}_{i}^{(r)}=\bm{W}^{V}\bm{h}_{i}^{(r-1)},$ (34)
@@ V-B1 Multi-Head Attention (MHA) layer
Based on (32), (33), and (34), we compute the compatibility $\bm{u}_{ij}^{(r)}\in\mathbb{R}$ by combining the query $\bm{q}_{i}^{(r)}$ with the key $\bm{k}_{j}^{(r)}$ of miner $j$ as the dot-product function[31]
$\bm{u}_{ij}^{(r)}=\begin{cases}\frac{\left(\bm{q}_{i}^{(r)}\right)^{T}\bm{k}_{j}^{(r)}}{\sqrt{d_{k}}}&\text{if miner $i$ is adjacent to miner $j$,}\\ -\infty&\text{otherwise,}\end{cases}$ (35)
where the compatibility of non-adjacent miners is considered as $\infty$- ∞, which is effective in preventing block propagation between non-adjacent miners. Then, based on the compatibilities $\bm{u}_{ij}^{(r)}$, the attention weight $\bm{a}_{ij}^{(r)}\in[0,1]$ is given by[25]
$\bm{a}_{ij}^{(r)}=softmax\left(\bm{u}_{ij}^{(r)}\right)=\frac{e^{\bm{u}_{ij}^{(r)}}}{\sum_{j=1}^{J}e^{\bm{u}_{ij}^{(r)}}},$ (36)
where $\{1,\ldots,j,\ldots,J\}\subset\mathcal{M}$ is the adjacent miner set of miner $i$. Then, based on (34) and (36), the result vector $\bm{h}_{i}^{\prime}(r)$, combining $\bm{a}_{ij}^{(r)}$ with $\bm{v}_{j}^{(r)}$, is expressed as[25]
$\bm{h}_{i}^{\prime}{}^{(r)}=\sum_{j=1}^{J}\bm{a}_{ij}^{(r)}\bm{v}_{j}^{(r)},$ (37)
@@
V-B2 Batch Normalization (BN) layer
@@
V-B3 Feed-Forward (FF) layer
@@
V-B4 Graph attention network layer
@@ V-C Decoder
Here, [,,,][$\cdot$,$\cdot$,$\cdot$,][ ⋅ , ⋅ , ⋅ , ] is the horizontal concatenation operator. For each time $t$, we utilize the graph embedding $\bar{\bm{h}}^{(graph)}$ as a part of the context vector $\bm{h}^{(d)}$, for the reason that the graph embedding is designed for taking the complete graph structure into account. For $t=1$, we introduce learnable $d_{h}$-dimensional parameters $\bm{v}^{1}$ and $\bm{v}^{2}$ as input placeholders. For $t=2$, the placeholder $\bm{v}_{1}$ would be replaced by the embedding $\bm{h}_{\pi_{1}}^{(N)}$. For $t>2$, the placeholder $\bm{v}_{2}$ and the embedding $\bm{h}_{\pi_{1}}^{(N)}$ would be replaced by the embeddings $\bm{h}_{\pi_{t-1}}^{(N)}$ and $\bm{h}_{\pi_{t-2}}^{(N)}$ of miners $\pi_{t-1}$ and $\pi_{t-2}$, to achieve the local optimum of AoB, and eventually achieve the optimization objective of minimizing overall AoB. Note that the result vector $\bm{h}^{(d)}$ is $(3\cdot d_{h})$-dimension to align with the miner embeddings $\bm{h}_{\pi_{t}}^{(N)}$.
1 for ${i}=1,\ldots,M$ do
2       Record the miner interactions $\alpha^{t_{k}}$ and $\beta^{t_{k}}$.
3      Calculate the miner reputation $R_{i\to j}^{final}$.
4
5Initialize $\bm{\theta}$, $\bm{\theta}^{BL}$leftarrow $\bm{\theta}$.
6for $epoch=1,\ldots,E$ do
7       for $step=1,\ldots,T$ do
8             $x_{i}\leftarrow$ RandomInstance( )[25], $\forall{i}\in\{1,\ldots,B\}$.
9            for ${i}=1,\ldots,M$ do
10                   if $R_{i\to j}^{final}$ < $\sigma$ then
11                         Mask miner $i$ to avoid the low-reputation miners.
12
13
14            $\bm{\pi}_{i}\leftarrow$ SampleRollout($G_{i},\bm{p}_{\bm{\theta}}$)[25], $\forall{i}\in\{1,\ldots,B\}$.
15            $\bm{\pi}_{i}^{BL}\leftarrow$ GreedyRollout($G_{i},p_{\bm{\theta}}^{BL}$)[25], $\forall{i}\in\{1,\ldots,B\}$.
16            $\nabla\mathcal{L}\leftarrow\sum_{i=1}^{B}\left(L(\bm{\pi_{i}|G)})-L(\bm{\pi_{i}^{\mathrm{BL}}|G)})\right)\nabla_{\bm{\theta}}\log p_{\bm{\theta}}(\bm{\pi_{i}})$.
17            $\bm{\theta}\leftarrow$ Adam($\bm{\theta},\nabla\mathcal{L}$).
18
19      if OneSidedPairedTTest($p_{\bm{\theta}},p_{\bm{\theta}}^{BL}$) < $\alpha^{gat}$[25] then
20             $\bm{\theta}^{BL}\leftarrow\bm{\theta}$.
21
22
23Feed the miners with their reputations into the trained network.
24Calculate the minimized overall AoB and the total value of miner reputations.
Output: The optimal block propagation trajectory, the minimized overall AoB, and the total value of miner reputations.
@@
(a) Channel bandwidth $b=180kHz$.
(b) Channel bandwidth $b=22MHz$.
(c) Channel bandwidth $b=100MHz$.
Figure 3: The overall AoB of block propagation corresponding to different numbers of miners in different channel bandwidths.
@@ VI Model Training
Furthermore, concerning the algorithm itself, each rollout incurs an additional forward pass, resulting in a $50\%$ increase in computational requirements[25]. However, given that the baseline policy remains constant throughout an entire epoch, we can streamline the process of sampling data and computing baselines by employing larger batch sizes. This efficiency is possible due to reduced memory requirements, allowing computations to operate in pure inference mode [25]. Note that the computational complexity of Algorithm 1 is $O(M)$[37], indicating that Algorithm 1 is efficient. Finally, after training the network, the miners, along with their reputations, are input into the trained network. The network minimizes the overall AoB of block propagation and calculates the total reputation of miners, generating the optimal block propagation trajectory for a specific number of miners as the final output.
VII Numerical Results
@@
In this section, we first compare the proposed GAT model with other conventional routing mechanisms: i) Greedy mechanism. In the simulation settings, the greedy mechanism always calculates the distance between miners and consistently selects the nearest adjacent miner to propagate the block; ii) Gossip mechanism. In the simulation settings, the gossip mechanism only takes into account adjacent miners, and the miner randomly propagates the block to their adjacent miner. At the same time, we separately add the reputation constraints to the three proposed mechanisms. Then, we present the optimal block propagation trajectories under different miner numbers in the proposed GAT model. Finally, we evaluate the impact of baselines and learning rate settings on the performance of the proposed GAT model.
@@
For the parameter settings of the experiment, the parameters $\bm{\theta}$ of the GAT network are initialized uniformly within the range $\big{(}-1/\sqrt{d_{f}},1/\sqrt{d_{f}}\big{)}$, where $d_{f}=2$ represents the input dimension of miners. Moreover, we use $N=3$ layers in the encoder, which is a thoughtful balance between the quality of results and computational complexity. We set $r^{ratio}=3/4$ to ensure that the block propagates to miners over $50\%$ for validation. Additionally, the training spans $100$ epochs using dynamically generated training data, processing $2500$ batches of $512$ instances per epoch. Note that we run the experiments on NVIDIA GeForce RTX 3080Ti, and the main parameters of this paper are listed in Table II[38, 39, 40, 11, 41].
@@
Fig. 3 shows the overall AoB of block propagation corresponding to different numbers of miners in different channel bandwidths utilizing different routing mechanisms. For convenience, we denote "GAT with Reputation Mechanism" to represent the GAT model with the reputation mechanism, which is the same as the "Greedy with Reputation Mechanism" and "Gossip with Reputation Mechanism". Considering different communication techniques, we conduct simulations with bandwidths of $b=180$$\rm{kHz}$[38], $b=22$$\rm{MHz}$[39], and $b=100$$\rm{MHz}$[40], corresponding to typical channels, WiFi, and 5G, respectively. From Fig. 3, we can observe that, as the channel bandwidth increases, the AoB decreases significantly, highlighting the necessity of improving channel bandwidth. Moreover, we can observe that both the proposed GAT model and the GAT model with the reputation mechanism consistently outperform the other four mechanisms. When $M=49$, the advantage is particularly pronounced, where the AoB of the GAT model is $2.5\%$ lower than that of the Greedy mechanism and $24.4\%$ lower than that of the Gossip mechanism, while the AoB of the GAT model with the reputation mechanism is $2.4\%$ lower than the Greedy mechanism with the reputation mechanism and $22.6\%$ lower than that of the Gossip mechanism with the reputation mechanism. The reason is that the GAT model focuses on minimizing the overall AoB of block propagation by considering the global structure of the miner network, which can obtain the optimal block propagation trajectory, thereby achieving block propagation optimization. In contrast, the Greedy mechanism only focuses on the selection of the current step and ignores the global structure, while the Gossip mechanism randomly selects miners without optimizing AoB, resulting in a much lower effect than other mechanisms.
Fig. 4 depicts the total value of miner reputations corresponding to different numbers of miners under different routing mechanisms, displaying a monotonic rise with an increase in the number of miners. As we can see, the algorithms with the miner reputation mechanism can obtain the propagation trajectory with a high total of value miner reputations, thus effectively ensuring reliability in block propagation. Moreover, the GAT model with the reputation mechanism performs best among other mechanisms and the reputation gap becomes more obvious with a larger number of miners. For example, comparing the Greedy mechanism with the GAT model with the reputation mechanism, the reputation of the GAT model with the reputation mechanism is $7.4\%$ higher than that of the Greedy mechanism at $M=9$, which outstandingly increases to $34.77\%$ at $M=49$. Combining Fig. 3 and Fig. 4, we can conclude that algorithms with the reputation mechanism inevitably improve the reliability of block propagation at the cost of increasing AoB [10]. However, the improvement in the reliability of block propagation with a relatively small increase in AoB is reasonable. Overall, the GAT model with the reputation mechanism stands out as the most exceptional performer among other mechanisms.
In Fig. 5, we present the optimal block propagation trajectory corresponding to different miner numbers $M=\{9,19,29,39,49\}$. For better distinction, we draw yellow points to denote the low-reputation miners and green points to denote the high-reputation miners, where the reputation values are calculated by the proposed reputation mechanism[30]. Moreover, the blue arrows point to the optimal block propagation trajectories generated by the proposed GAT model. We can observe that the proposed GAT model can obtain optimal block propagation trajectories based on different numbers of miners, which are clearly organized and have no unreasonable costs in AoB for efficiency. At the same time, the block propagation trajectories can perfectly prevent the blocks from propagating to the low-reputation miners, which ensures the reliability of block propagation.
Fig. 6 depicts the network loss corresponding to training epochs with four learning rate schemes for miner numbers $M=19$ and $M=49$. As shown in Fig. 6(a), the GAT model with the learning rate $lr=10^{-3}$ outperforms the GAT model with the learning rate $lr=10^{-4}$, with marginal improvement from learning rate decay[25]. However, in Fig. 6(b), the loss of the GAT model with the learning rate $lr=10^{-3}$ converges slowly and fails to fully converge after $100$ epochs with a larger loss. Conversely, employing a learning rate of $lr=10^{-3}$ with a decay of $0.96^{\text{epoch}}$ yields better model performance, and further improvement may be achievable with prolonged training. Furthermore, motivated by the above analysis, we conclude that employing a higher learning rate of $lr=10^{-3}$ accelerates the initial learning but necessitates decay for convergence[25], especially in the case of a larger number of miners. Moreover, a smaller learning rate of $lr=10^{-4}$ shows increased stability without the need for decay, albeit with a limited impact on loss reduction.